OpenStudy (anonymous):

Choose the correct simplification of the expression (a^2)^3. a^5 a^8 a a^6

OpenStudy (anonymous):

\[\huge (a^b)^c = a^{b \times c}\]

OpenStudy (anonymous):

\[(b^m)^n=b^{m\times n}\]

OpenStudy (anonymous):

It the answer A. or D.

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

@satellite73